Tswer: D. 36) One mole of potassium sulfate contains:
60) Determine the correct empirical formula of a compound containing 26.68% carbon, 2.24% hydrogen, and 71,08% oxygen.
Answer #1

Answer:-

(36) -

As we know that a mole is defined as the amount of substance containing chemical particles like atoms , molecules or ions equal to the Avogadro's number i.e 6.023 \times 1023 e.g.

1 mole Al = 27 g Al = 6.023 \times 1023 Al atoms

1 mole H2O = 18 g H2O = 6.023 \times 1023 H2O molecules

1 mole NaCl  = 58.5 g H2O = 6.023 \times 1023 NaCl molecules

therefore

1 mole potassium sulfate contains = 1 mole K2SO4  = 174.26 g K2SO4 = 6.023 \times 1023 K2SO4 molecules  (i.e the answer)

(60) -

Given:-

percentage of carbon (C) = 26.68 %

percentage of hydrogen (H) = 2.24 %

percentage of oxygen (O) = 71.08 %

As we know that empirical formula is the formula of compound which gives the simple whole number ratio of atoms of various elements present in one molecule of the compound.

therefore following table in which simple whole number ratio of carbon (C) , hydrogen (H) and oxygen (O) present in the compound can be calculated as follows:-

Element Symbol percent by mass (%) Atomic mass Moles of element (% / At. mass) Simple molar ratio Simplest whole no. molar ratio
Carbon C 26.68 12 26.68 / 12 = 2.22 2.22 / 2.22 = 1 1
Hydrogen H 2.24 1 2.24 / 1 = 2.24 2.24 /2.22 = 1 1
Oxygen O 71.08 16 71.08 / 16 = 4.44 4.44 / 2.22 = 2 2

So

correct empirical formula of compound = C1H1O2

or

correct empirical formula of compound = CHO2 (i.e the answer)